OpenStudy (anonymous):

Use the iterative rule to find the 8th term in the sequence. an = 25 – 3n

OpenStudy (mathmath333):

\(\Large\tt \begin{align} \color{black}{a_n = 25 – 3n\\ a_8 = 25 – 3\times 8\\ ~~~~=25-24\\ ~~~~=1}\end{align}\)